OpenStudy (anonymous):

Find all points on the curve x^2 * y^2 + xy = 2 where the slope of the tangent line is -1.

OpenStudy (anonymous):

I got the derivative already: y' = (-2xy^2 - y)/(2x^2 * y + x) ...=-1 so how do I solve for those points?

OpenStudy (anonymous):

right coulda guessed that much and

OpenStudy (inkyvoyd):

I think you substitute -1 for all values of y.

OpenStudy (inkyvoyd):

Then you solve. I think. Check your answers afterwards to see it I was right.

OpenStudy (inkyvoyd):

*y' I mean

OpenStudy (inkyvoyd):

for all values of y', we should have -1

OpenStudy (anonymous):

right

OpenStudy (inkyvoyd):

wait, I figured it out for a circle.

OpenStudy (inkyvoyd):

x^2+y^2=sqrt(2) 2x+2yy'=0 y'=-1 because the slope=-1 2x-2y=0 x=y so we have to equations x=y x^2+y^2=sqrt2

OpenStudy (inkyvoyd):